### v5

# Toward expert-level medical text validation with language models

Abstract
With the growing use of language models (LMs) in clinical environments, there is an immediate need to evaluate the accuracy and safety of LM-generated medical text. Currently, such evaluation relies solely on manual physician review. However, detecting errors in LM-generated text is challenging because 1) manual review is costly and 2) expert-composed reference outputs are often unavailable in real-world settings. While the “LLM-as-a-judge” paradigm (a LM evaluating another LM) offers scalable evaluation, even frontier LMs can miss subtle but clinically significant errors. To address these challenges, we propose MedVAL, a novel, self-supervised, data-efficient distillation method that leverages synthetic data to train evaluator LMs to assess whether LM-generated medical outputs are factually consistent with inputs, without requiring physician labels or reference outputs. To evaluate LM performance, we introduce MedVAL-Bench, a dataset of 840 physician-annotated outputs across 6 diverse medical tasks capturing real-world challenges. Each output is reviewed following a physician-defined taxonomy of risk levels and error categories, enabling evaluation of LMs in making deployment safety decisions. Across 10 state-of-the-art LMs spanning open-source, proprietary, and medically adapted models, MedVAL distillation significantly improves ($p<0.001$) alignment with physicians across seen and unseen tasks, increasing average F1 scores from 66% to 83%. Despite strong baseline performance, MedVAL improves the best-performing proprietary LM (GPT-4o) by 8% without training on physician-labeled data, demonstrating a performance statistically non-inferior to a single human expert on a subset annotated by multiple physicians ($p<0.001$). To support a scalable, risk-aware pathway towards clinical integration, we open-source: 1) Codebase, 2) MedVAL-Bench, 3) MedVAL-4B. Our benchmark provides evidence of LMs approaching expert-level ability in validating AI-generated medical text.
1 Introduction
Figure 1: a) MedVAL test-time workflow. A generator LM produces an output, and MedVAL then assesses the output’s factual consistency with the input, while assigning a risk grade and determining whether the output is safe for deployment or not. b) Study framework. 12 physicians assess 840 LM-generated medical text outputs. Using physician assessments as reference, we measure the accuracy of LMs in medical text validation across 10 LMs, 2 methods (baseline vs. MedVAL), and 6 tasks.
Language models (LMs) are increasingly leveraged for documentation in medical settings, supporting tasks such as clinical text summarization, report generation, or question answering [1, 2, 3, 4, 5, 6, 7]. While LMs offer potential for reducing documentation burden [8, 9, 10, 11], they are not infallible and frequently generate plausible text that contains subtle errors such as hallucinations, omissions, or certainty misalignments [12, 13, 14]. This issue is critical within medicine, where errors are often cloaked in jargon, misleading even experienced practitioners [15].
The adoption of LMs [16, 17, 18] for medical applications necessitates reliable risk assessment through medical text validation, which involves determining whether an output from a LM is factually consistent with the input. Currently, this process requires manual physician review both pre- and post-deployment to ensure continued safety, contributing to documentation workload and associated cognitive fatigue for physicians [19, 20, 21, 22].
Despite this need, automatically evaluating LM-generated outputs remains difficult to scale. Traditional NLP metrics fail to capture subtle, clinically significant errors that can impact patient care [23]. These metrics rely on expert-composed outputs (reference outputs), which are not available in many real-world scenarios [24]. Moreover, manual physician review remains the reference standard for validating medical text, but it is costly, time-consuming, and not scalable. With physician shortages and burnout at an all-time high [25, 26, 27, 28], assigning physicians additional validation tasks is impractical. Therefore, automated and scalable strategies that minimize reliance on expert-driven data are urgently needed to adapt LMs for medical text validation.
Using a LM to evaluate another LM’s output, also termed the "LLM-as-a-judge" paradigm, has shown promise in automating the evaluation of LM-generated text. However, most general-purpose LLM-as-a-judge methods treat LMs as static evaluators [29], often lacking the granularity required to assess nuanced, high-stakes medical text [30, 31, 32, 33, 34]. Recent medical-specific LLM-as-a-judge approaches suffer from certain practical limitations; these methods often assume resources like expert-labeled training data [35], availability of reference outputs [36], or retrieval-based evidence [37], none of which are readily scalable. On the other hand, many approaches focus on subdomains such as radiology (e.g., chest X-rays), limiting generalization to broader medical tasks [38, 39, 40, 41, 42, 43, 44].
To overcome these challenges, we introduce MedVAL (medical text validator), a self-supervised, data-efficient distillation method (Figure 1), which curates high-quality synthetic examples by leveraging the agreement between a generator and a validator LM as a proxy for physician judgment [45]. Through this training, MedVAL enables LMs to assess whether an output is factually consistent with the input by assigning one of four risk levels, while flagging "deployment unsafe" outputs at near physician-level reliability. In comparison to prior methods, our framework enables: 1) development of scalable evaluators without physician-in-loop supervision, 2) evaluation without reliance on reference outputs, and 3) adaptation across diverse tasks.
We evaluate MedVAL across 6 open-source and 4 proprietary LMs on MedVAL-Bench [46], our newly created dataset consisting of 840 physician-annotated outputs, where each output is annotated following a physician-defined taxonomy of risk levels and error categories. MedVAL-Bench includes 6 diverse tasks (focused on LM output evaluation), including medication question answering, patient query summarization, radiology report summarization, radiology impression simplification, hospital course translation, and doctor-patient dialogue summarization. To reflect the multilingual nature of U.S. clinical practice, where Spanish is the second most spoken language [47], we include an English-to-Spanish translation task annotated by bilingual physicians.
Unlike prior LLM-as-a-judge methods that rely on zero-shot inference, MedVAL improves the validation capabilities of all underlying LMs across seen and unseen tasks. MedVAL yields significant gains ($p<0.001$): average baseline F1 scores for four-class risk-grading improve from 36.7% to 51.0%, and for binary safe/unsafe classification, from 66.2% to 82.8%. Our benchmark provides evidence that LMs can achieve performance statistically non-inferior to a single human expert on a subset annotated by multiple physicians ($p<0.001$).
Figure 2: MedVAL self-supervised data curation illustrated through a radiology report summarization example. 1) A generator $g_{\theta}$ takes as input $x$, and generates a clean and a perturbed output using a random perturbation level $\delta\in[0,1]$. 2) A validator $v_{\phi}$ then provides a detailed error assessment, predicts the factual degradation level $\hat{\delta}_{clean}$ and $\hat{\delta}_{corrupt}$ of the clean and perturbed outputs, respectively, and filters data with high generator-validator consistency for fine-tuning an arbitrary LM.
2 Methods
2.1 MedVAL Training
We propose a self-supervised distillation method for training LMs to validate LM-generated medical text, involving three stages: 1) synthetic data generation, 2) data filtering, and 3) fine-tuning. Figure 2 describes the data curation process, and Algorithm S1 summarizes the complete training process.
Definitions.
[…]
Stage 1: Synthetic data generation.
Our goal is to create a synthetic dataset $\mathcal{D}=\{x,\hat{y},\hat{y}_{\delta}\}$ that contains a triplet of the input, clean unperturbed output, and the perturbed output, to enable supervision without expert labels. To achieve this, we first introduce perturbations in the outputs. The perturbations in $\hat{y}_{\delta}$ are introduced via additional instructions in the generator’s prompt. To ensure each perturbation level corresponds to an interpretable perturbation method, we define a discrete set of degradation levels $\{\delta_{1},\delta_{2},\ldots,\delta_{L}\}$. These degradation levels were categorized into four clinically meaningful strata—no risk, low risk, moderate risk, and high risk—in conjunction with our physician team, based on their potential to influence patient safety. While we use a discrete risk-level taxonomy for implementation, the framework supports continuous degradation levels and scalar validator outputs, which can be modified based on the perturbation injection method. For each sample in the dataset, we randomly select a perturbation level to ensure coverage of a range of error severities. The perturbations, associated risk levels, and instructional prompts are described in Table 1.
Table 1: Physician-defined taxonomy of risk levels, safety rules, perturbation instructions, and error categories.
Perturbation Category Risk Safety Action
$\delta=0\%$ Level 1 No Risk Safe Expert review not required.
$\delta=33\%$ Level 2 Low Risk Acceptable Expert review optional.
$\delta=67\%$ Level 3 Moderate Risk Potentially unsafe Expert review required.
$\delta=100\%$ Level 4 High Risk Unsafe Expert rewrite required.
Perturbation Instructional prompt
$\delta=0\%$ "The output should contain no clinically meaningful factual inconsistencies. Any deviations from the input (if present) should not affect clinical understanding, decision-making, or safety."
$\delta=33\%$ "The output should contain subtle or ambiguous inconsistencies that are unlikely to influence clinical decisions or understanding. These inconsistencies should not introduce confusion or risk."
$\delta=67\%$ "The output should contain inconsistencies that could plausibly affect clinical interpretation, documentation, or decision-making. These inconsistencies may lead to confusion or reduced trust, even if they don’t cause harm."
$\delta=100\%$ "The output should include one or more inconsistencies that could result in incorrect or unsafe clinical decisions. These errors should pose a high likelihood of compromising clinical understanding or patient safety if not corrected."
Error category Error Description
Hallucinations Fabricated claim Introduction of a claim not present in the input.
Misleading justification Incorrect reasoning, leading to misleading conclusions.
Detail misidentification Incorrect reference to a detail in the input.
False comparison Mentioning a comparison not supported by the input.
Incorrect recommendation Suggesting a diagnosis/follow-up outside the input.
Omissions Missing claim Failure to mention a claim present in the input.
Missing comparison Omitting a comparison that details change over time.
Missing context Omitting details necessary for claim interpretation.
Certainty Misalignments Overstating intensity Exaggerating urgency, severity, or confidence.
Understating intensity Understating urgency, severity, or confidence.
Other Other Additional errors not covered.
Stage 2: Data filtering.
We aim to filter $\mathcal{D}$ to obtain a "high-quality" subset $\mathcal{D}_{\text{train}}$ for fine-tuning a validator LM. $\mathcal{D}_{\text{train}}$ should ideally contain examples where the validator $v_{\phi}$ and generator $g_{\theta}$ strongly agree on the expected factual degradation level. To achieve this, we first leverage the validator $v_{\phi}$ to predict the factual degradation levels for both the clean and perturbed outputs. For filtering the synthetic data as a "quality control" mechanism, we propose a metric $\mathcal{M}_{\text{MedVAL}}$ that quantifies consistency between the generator and validator via evaluating the agreement between expected and predicted factual consistency levels. We propose two independent components for our metric. 1) Absolute consistency ensures the validator’s predicted factual degradation level for each generation should match the expected degradation level. Specifically, for clean generations, we expect $v_{\phi}(x,\hat{y})\approx 0$. For perturbed generations, we expect $v_{\phi}(x,\hat{y}_{\delta})\approx\delta$. 2) Relative consistency ensures the validator preserves the expected factual degradation between clean and perturbed generations. Specifically, we expect $v_{\phi}(x,\hat{y}_{\delta})-v_{\phi}(x,\hat{y})\approx\delta$. We define these terms as:
$\mathcal{M}_{\text{consistency}}=\underbrace{\|v_{\phi}(x,\hat{y})\|_{2}^{2}+\|v_{\phi}(x,\hat{y}_{\delta})-\delta\|_{2}^{2}}_{\mathcal{M}_{\text{absolute}}}+\underbrace{\|v_{\phi}(x,\hat{y}_{\delta})-v_{\phi}(x,\hat{y})-\delta\|_{2}^{2}}_{\mathcal{M}_{\text{relative}}}$ (1)
The first term in $\mathcal{M}_{\text{absolute}}$ rewards clean generations that achieve validator scores close to 0, and the second term rewards perturbed generations whose validator scores closely match the expected degradation $\delta$. $\mathcal{M}_{\text{relative}}$ ensures that the difference between the predicted scores reflects the expected degradation. To bound the score between 0 and 1 ($\uparrow$ score == $\uparrow$ generator-validator consistency), we divide it by its maximum possible value:
$\mathcal{M}_{\text{MedVAL}}=1-\frac{\mathcal{M}_{\text{consistency}}}{6}$ (2)
Stage 3: Fine-tuning.
We create our synthetic dataset $\mathcal{D}_{\text{train}}=\{x,\hat{y},\hat{\delta}_{\text{clean}},\hat{y}_{\delta},\hat{\delta}_{\text{corrupt}}\}$ by filtering examples from $\mathcal{D}$ where $\mathcal{M}_{\text{MedVAL}}\geq\tau$ (pre-specified threshold). Importantly, we apply a single-pass consistency filter (not iterative), and choose $\tau$ to retain examples with high generator-validator agreement. We then fine-tune an arbitrary LM $v^{*}_{\alpha}=\textit{SFT}(v_{\alpha},\mathcal{D}_{\text{train}})$ using standard parameter-efficient supervised fine-tuning (SFT). To improve robustness, we optionally ensemble multiple MedVAL fine-tuned LMs by aggregating their predictions.
Table 2: MedVAL-Bench dataset. Overview of task-specific data sources, relevant statistics, and generation instructions. In-dist denotes tasks seen during LM fine-tuning. # physicians denotes number of physicians involved in test-time annotation. The train set is self-supervised (no physician labels) and relies on input-output pairs to create synthetic supervision labels.
In- # samples Avg. # tokens # phys-
Task Dataset dist Train Test Train Test icians
medication2answer MedicationQA 500 135 10 $\pm$ 4 10 $\pm$ 4 2
query2question MeQSum 500 120 80 $\pm$ 57 82 $\pm$ 66 3
report2impression Open-i 500 190 54 $\pm$ 21 50 $\pm$ 22 5
report2simplified Open-i 500 - 54 $\pm$ 22 - -
impression2simplified MIMIC-IV - 190 - 69 $\pm$ 61 5
bhc2spanish MIMIC-IV-BHC - 120 - 543 $\pm$ 391 3
dialogue2note ACI-Bench - 85 - 1,497 $\pm$ 445 2
Task Input $\to$ output Instructional prompt
medication2answer medication question $\rightarrow$ answer “Answer the following medication-related
patient health question.”
query2question patient query $\rightarrow$ health question “Summarize the patient health query into
one question of 15 words or less.”
report2impression findings $\rightarrow$ impression “Summarize the radiology report findings
into an impression with minimal text.”
report2simplified findings $\rightarrow$ patient-friendly “Create a simplified, patient-friendly
version of the input.”
impression2simplified impression $\rightarrow$ patient-friendly “Create a simplified, patient-friendly
version of the input.”
bhc2spanish hospital course $\rightarrow$ spanish “Translate the brief hospital course into
Spanish.”
dialogue2note doctor-patient dialogue $\rightarrow$ note “Summarize the doctor/patient dialogue
into an assessment and plan.”
2.2 Data
We introduce MedVAL-Bench (Table 2), a dataset for evaluating the ability of LMs to validate LM-generated medical text. The train set is self-supervised, containing inputs (used to generate supervision labels) and reference outputs (if available). The test set additionally contains gold-standard physician assessments of outputs. Next, we describe: 1) train/test input sources, and 2) test-time data curation via the reader study.
2.2.1 Input Sources
Medication question answering.
The medication2answer task involves answering a medication-related question, and utilizes the MedicationQA dataset [48], which contains real-world consumer health questions about medications and expert-written answers. Unlike other tasks, the input (a standalone question) often lacks sufficient context to validate the output’s factual consistency. Hence, we aim to evaluate whether LMs can reliably perform safety-critical decisions using their knowledge base and the original question. Due to the unavailability of pre-specified train/test splits, we randomly create the splits as specified in Table 2.
Patient query summarization.
The query2question task involves creating a concise query that encapsulates the essential details necessary to address the original question. We utilize the MeQSum dataset [49], which comprises: 1) health-related queries from patients, sourced from messages sent to the U.S. National Library of Medicine, and 2) corresponding abbreviated questions formulated by three medical experts to ensure that the summaries facilitate complete and accurate answer retrieval. For our study, we sample a subset of examples from both the train and test splits defined by Van Veen et al. [4], as detailed in Table 2.
Radiology report summarization.
The report2impression task involves processing the "findings" section of a radiology exam report to create an "impression" section that succinctly highlights the critical information. We leverage the Open-i dataset [50], which consists of de-identified narrative chest x-ray reports from the Indiana Network for Patient Care database. We randomly sample train/test examples following the splits identified by Van Veen et al. [4], ensuring the number of train/test samples indicated in Table 2.
Radiology impression simplification.
The impression2simplified task involves simplifying radiology report impression sections to produce patient-friendly rewrites that preserve the clinical meaning while being understandable to a layperson. For this task, we extract the "impression" sections from the MIMIC-IV dataset [51], which contains comprehensive radiology reports sourced from patient stays in critical care units of the Beth Israel Deaconess Medical Center. Our sampling process ensured an equal distribution of imaging modalities that reflect the top imaging indications at Stanford Hospital. These include chest x-rays, CT scans of the abdomen and pelvis, CT scans of the head, MR brain studies, ultrasound exams of the pelvis, digital screening mammography, and transabdominal/transvaginal pelvic ultrasounds. Test examples were selected via stratified random sampling as indicated in Table 2. Because MIMIC-IV [51] is partially open-source, we select training examples from Open-i radiology report findings on a related but distinct report2simplified task (i.e., simplifying the findings section). For this task, we randomly sample training examples from the training split by Van Veen et al. [4], ensuring no overlap with the samples used in report2impression.
Hospital course translation.
The bhc2spanish task involves translating brief hospital course (BHC) sections from discharge summaries from English into Spanish. We employ the MIMIC-IV-BHC dataset [52], which contains curated BHC sections from discharge summaries, written by healthcare providers at the Beth Israel Deaconess Medical Center. The MIMIC-IV-BHC dataset includes diverse patient encounters, ensuring a comprehensive representation. We randomly sample testing examples as indicated in Table 2.
Doctor-patient dialogue summarization.
The objective of dialogue2note is to summarize doctor-patient conversations into an "assessment and plan". We utilize the ACI-Bench dataset [53, 54, 55], which provides a collection of 1) 207 doctor-patient conversations and 2) patient visit notes. To ensure consistency, we randomly sample the testing examples (indicated in Table 2), from the test set defined by Van Veen et al. [4].
2.2.2 Physician Reader Study
We design our reader study to enable a robust evaluation of how well LMs assess the factual consistency of outputs, compared to physicians. For each test-time input, an LM-generated output was produced by sampling a perturbation level $\delta\in[{0,0.33,0.67,1.0}]$ uniformly at random, and injecting the corresponding perturbation instruction into the prompt as described in Table 1. This ensured that the test set contained a balanced mix of outputs with no risk, low risk, moderate risk, and high risk errors. Since our objective was to evaluate the validator’s ability to distinguish clinically meaningful degradation, we do not explicitly prompt the generator to produce only no-risk outputs, as doing so would instead assess generation quality. This design was necessary to evaluate a validator’s ability to distinguish degradation across a meaningful severity spectrum. We observe that models sometimes diverge from requested perturbations; consequently, outputs often contain a blend of prompted and natural errors committed by LMs. In all cases, physician graders assign risk levels based solely on the observed text, independent of the prompt specification. Importantly, these perturbation rules apply to the test set, as the training set is curated using the pipeline in Figure 2.
To curate reference evaluations of LM outputs, we assembled a panel of 12 physicians, where 4 board-certified internal medicine physicians annotated general medical tasks (medication2answer, query2question, dialogue2note), 3 bilingual internal medicine residents annotated the bhc2spanish task, and 4 board-certified radiologists and a radiology resident annotated radiology tasks (report2impression, impression2simplified). For each study, physicians were presented with the input and LM-generated output (without indicating error injection), and requested to perform: 1) risk grading: "assign a risk level to the output, following the risk level taxonomy (between 1 and 4)", and 2) error assessment: "identify factual consistency errors in the output based on the error category taxonomy (hallucinations, omissions, or certainty misalignments)". 15 random examples from each task were annotated by multiple physicians, allowing us to measure inter-physician agreement.
We present the risk and error category distribution in Table S3 and Figure S6. The most common error categories are fabricated claim (45.7%), missing claim (14.0%), and incorrect recommendation (12.6%). We observe that error frequency rises with the risk grades (for grades 1–4, average 0.14 $\to$ 3.24 errors).
2.3 Experimental Setup
2.3.1 Language Models
We evaluate a diverse collection of state-of-the-art transformer-based LMs (Table S2). Our selection criteria included model license, context length, and size. For open-source models, we include Llama [56, 57] (Llama 3.1 8B, 3.2 3B, and 3.3 70B), Qwen3 [58] (4B dense version), and Gemma [59] (Gemma 3 27B and MedGemma 27B). For proprietary models, we include GPT-4o Mini, GPT-4o [60, 61], Claude Sonnet 4, and Gemini 2.0 Flash [62]. The proprietary models directly serve as our state-of-the-art LLM-as-a-judge baselines [63].
2.3.2 Implementation Details
Under open-source category, we select LMs under 8 billion parameters for fine-tuning, as they can be easily deployed on consumer-grade GPUs. We leverage DSPy’s bootstrap fine-tuning algorithm [64, 65]. As part of our contribution, we extend DSPy’s existing local parameter-efficient fine-tuning pipeline to enable Quantized Low-Rank Adaptation (QLoRA) (GitHub PR) [66]. We employ an NVIDIA A6000 GPU and fine-tune models for 5 epochs using the Adam optimizer [67]. We initialize the learning rate at $1\times 10^{-5}$ with linear decay. We set the per-device train batch size to 1 and apply 4-bit precision quantization using BitsAndBytesConfig [68].
Under proprietary LMs, we report zero-shot baseline performance and only fine-tune OpenAI models that allow fine-tuning, where DSPy collects training data and initiates a fine-tuning procedure managed by OpenAI. Since OpenAI handles the parameter-efficient fine-tuning internally, exact details are not publicly available.
Unless specified, our framework employs GPT-4o as the curator of synthetic data under stages 1 and 2 from Figure 2, and filters synthetic data using $\mathcal{M}_{\text{MedVAL}}\geq\tau$, where $\tau=0.9$ (selected heuristically to reflect high-consistency). All results are reported using DSPy’s zero-shot chain-of-thought (CoT) prompting module.
2.3.3 Statistical Analysis
We compute the F1 score for each risk level and task independently, comparing LM-predicted risk classifications to physician labels. The F1 score is a measure of a model’s accuracy that considers both precision (the correctness of the positive predictions) and recall (the ability to identify all positive instances). F1 score ranges from 0 to 1, where 1 indicates perfect precision and recall, and 0 indicates the worst performance with no correct positive predictions. We report F1 score using the macro average, ensuring equal weighting for risk levels and tasks to help mitigate class imbalance. For evaluating binary safety classification, we additionally report sensitivity (unsafe recall) and specificity (safe recall) to make class balance explicit.
We also report Cohen’s $\kappa$, allowing us to assess a LM’s agreement with physicians. Cohen’s $\kappa$ [69] is a statistical measure of agreement for categorical items, which takes into account the possibility of agreement occurring by chance. Given the ordinal nature of the four-class validation task, we compute Cohen’s $\kappa$ with linear weighting, comparing model-predicted risk levels to physician-assigned levels (1–4). The $\kappa$ score ranges from -1 to 1, where 1 represents perfect agreement, 0 indicates agreement no greater than chance, and -1 indicates complete disagreement. We use Krippendorff’s $\alpha$ [70] to assess inter-physician agreement over a subset of examples annotated by multiple physicians, evaluating the robustness of the physician annotations. Krippendorff’s $\alpha$ ranges from 0 to 1, where 1 indicates perfect agreement among annotators.
To assess whether MedVAL significantly improves performance over baseline models, we apply the McNemar test, a non-parametric test for paired nominal data that detects differences in classification outcomes. We evaluate whether MedVAL and the baseline model differ in their ability to correctly classify examples. To account for multiple comparisons (10 models $\times$ 4 risk levels), we apply a Bonferroni correction with a Type I error rate of $\alpha=0.05$. All statistical analyses were conducted using Python’s statsmodels package.
To test the non-inferiority (NI) of a LM against a single human expert, we use the 90/840 test cases with multiple-physician ratings. To identify a reference label, we use majority consensus. A "single expert" is defined by randomly sampling one physician’s label. We compute macro-F1 against the reference for both the model and sampled expert, form the NI contrast $\Delta$ = F1 (model) - F1 (single expert), and use a paired item-level bootstrap (B=10,000). Non-inferiority is declared if $\text{LCB}_{95\%}(\Delta)>-0.05$ (clinically negligible difference). We did not apply a multiplicity adjustment as our NI comparison was GPT-4o vs a single expert.
Figure 3: Performance benchmark (F1 score). a) We report the performance of LMs before and after MedVAL distillation. b) We rank all LMs (low to high), grouped into three methods. c) We report the $\Delta$ F1 score between MedVAL and baseline LM performance across each prediction class. Baseline indicates zero-shot LM before distillation, Baseline (larger comparator) indicates a larger zero-shot LM as reference (not chosen for distillation), and MedVAL indicates LM after distillation. $^{*}p<0.001$ indicates statistically significant difference in classification performance of MedVAL and baseline (McNemar test). Notably, smaller MedVAL LMs match or exceed the performance of much larger baseline LMs. Furthermore, MedVAL Qwen3-4B (52.7%) and GPT-4o (58.7%) achieve the highest F1 score ranking under respective categories.
3 Results
We compare two methods: 1) zero-shot baseline, and 2) MedVAL. Baseline involves directly prompting the LM without additional adaptation, whereas MedVAL refers to fine-tuning the model via MedVAL distillation.
3.1 Overall Performance
We show that MedVAL distillation consistently improves the performance of all LMs. The F1 score and Cohen’s $\kappa$, which reflect classification accuracy and agreement with expert-generated results, respectively, are reported in Figure 3 and Figure S1. Across all tested LMs, MedVAL distillation leads to average baseline F1 scores improving from 36.7% to 51.0%. On average, MedVAL improved F1 scores by 113% for open-source LMs and 11% for proprietary LMs, where baseline performance was already considerably high. Despite the inherent difficulty of this four-class classification task, where even frontier LMs struggle and baseline scores never exceed 55%, MedVAL distillation improves GPT-4o’s F1 score from 54.5% to 58.7%, establishing the best performance in our benchmark.
First, we examine performance across open-source models. Among all open-source LMs, Qwen-3-4B exhibits the highest F1 score after MedVAL distillation (42.8% $\to$ 52.7%), surpassing other notably larger open-source models such as Llama-3.3-70B and even the medically fine-tuned MedGemma-27B. Among zero-shot baselines, MedGemma-27B achieves the highest performance (48.2%), slightly outperforming Llama-3.3-70B and Gemma-3-27B. Notably, MedVAL distillation substantially improves the F1 scores of Llama-3.2-3B (12.8% $\to$ 42.9%), Llama-3.1-8B (25.9% $\to$ 46.5%). Despite Llama-3.1-8B also benefiting from MedVAL, MedVAL Qwen-3-4B outperforms it, emphasizing the advantage of specific baseline LM training techniques. Remarkably, MedVAL Qwen-3-4B even outshines certain proprietary baselines such as GPT-4o Mini (47.4%) and Gemini 2.0 Flash (51.5%). We refer to the MedVAL Qwen-3-4B model as “MedVAL-4B” for public release.
Within proprietary LMs, post MedVAL distillation, GPT-4o Mini (47.4% $\to$ 54.0%) rivals the baseline performance of larger Claude Sonnet 4 (54.2%) and GPT-4o (54.5%), and surpasses Gemini 2.0 Flash (51.5%), highlighting that MedVAL can enable LMs to reach or exceed baseline performance of much larger LMs.
We report Krippendorff’s $\alpha$ to assess inter-reader variability across the six tasks in Table 3a. The overall $\alpha$ of 0.848 indicates substantial to almost perfect agreement, proving the reliability of the physician annotations. A value of $\alpha\geq 0.80$ indicates reliable rating quality to draw triangulated conclusions based on the rated data [71]. Notably, the highest inter-physician agreement was observed in the bhc2spanish and medication2answer tasks, with an $\alpha$ of 0.943 and 0.904, respectively, while the lowest was in the query2question task, with an $\alpha$ of 0.560. We also observe that physician agreement correlates strongly with MedVAL performance: we observe a Pearson correlation of $r=0.67$ between inter-physician agreement and GPT-4o MedVAL F1 scores. This suggests that when experts agree more, MedVAL produces higher and consistent F1 scores.
3.2 Risk-Level Classification Performance
MedVAL improves classification performance across risk levels, demonstrating robustness. Figure 3c demonstrates the $\Delta$ F1 score between MedVAL and baseline LM performance. Gains are most pronounced in smaller open-source LMs, where improvements are statistically significant ($p<0.001$). For larger proprietary LMs, we also observe improvements, although they are not statistically significant ($p>0.1$).
Among open-source models, Llama-3.2-3B showed the largest gains, with F1 score increases of over 50% at levels 1 and 4, and over 17% at level 2. Qwen3-4B exhibited substantial gains, with improvements in level 2 (13.6%) and level 4 (20.2%). Llama-3.1-8B also improved substantially at level 1 (14.5%) and level 4 (55.2%).
Proprietary models like GPT-4o Mini exhibited notable gains at intermediate levels (level 2: 14.5%; level 3: 13.2%). Even GPT-4o, the strongest baseline model, showed smaller but consistent improvements, especially at level 2 (7.3%) and level 3 (6.9%). These gains are clinically significant because levels 2 and 3 represent the threshold between outputs that are likely safe for use with minimal oversight (level 2) and those requiring human review or revision (level 3). Interestingly, MedVAL sharpens this decision boundary for even high-performing baseline frontier LMs. Although minor drops were observed in a few model-risk combinations, such as GPT-4o Mini at level 4 (-2.2%) and Llama-3.2-3B at level 3 (-2.7%), the overall pattern shows strong and consistent gains across clinically meaningful risk categories.
Table 3: a) Task-wise performance benchmark (F1 score). Bolded values highlight best performance under respective categories (open-source/proprietary). Green and red values indicate improvement or decline in F1 score compared to the corresponding baseline LM. On average across LMs, MedVAL distillation increases F1 scores by 65% on in-distribution tasks and 84% on out-of-distribution tasks. Krippendorff’s $\alpha=0.848$ represents high inter-physician agreement across examples evaluated by multiple physicians, indicating the reliability of physician annotations. MedVAL Qwen3-4B (referred to as MedVAL-4B) achieves the best performance among all open-source LMs, matching the baseline performance of proprietary LMs.
In-Distribution Out-of-Distribution
Method Model medication2 answer query2 question report2 impression impression2 simplified bhc2 spanish dialogue2 note Overall
Open-Source
Baseline Llama-3.2-3B 0.091 0.110 0.174 0.096 0.120 0.146 0.128
Qwen3-4B 0.357 0.299 0.530 0.390 0.364 0.552 0.428
Llama3.1-8B 0.342 0.285 0.278 0.225 0.158 0.113 0.259
Gemma3-27B 0.398 0.279 0.584 0.442 0.369 0.552 0.459
MedGemma-27B 0.462 0.287 0.616 0.451 0.349 0.603 0.482
Llama-3.3-70B 0.478 0.311 0.633 0.496 0.362 0.322 0.480
MedVAL Llama-3.2-3B 0.382 +320% 0.262 +138% 0.578 +232% 0.429 +347% 0.242 +102% 0.448 +207% 0.429 +235%
Qwen3-4B 0.557 +56% 0.374 +25% 0.562 +6% 0.537 +38% 0.424 +16% 0.490 -11% 0.527 +23%
Llama-3.1-8B 0.456 +33% 0.372 +31% 0.480 +73% 0.540 +140% 0.384 +143% 0.376 +233% 0.465 +80%
Proprietary
Baseline GPT-4o Mini 0.479 0.352 0.445 0.503 0.427 0.586 0.474
GPT-4o 0.598 0.360 0.519 0.587 0.439 0.618 0.545
Claude Sonnet 4 0.569 0.413 0.497 0.583 0.552 0.550 0.542
Gemini 2.0 Flash 0.485 0.401 0.588 0.486 0.497 0.602 0.515
MedVAL GPT-4o Mini 0.512 +7% 0.308 -13% 0.635 +43% 0.571 +14% 0.386 -10% 0.692 +18% 0.540 +14%
GPT-4o 0.695 +16% 0.361 +0% 0.564 +9% 0.605 +3% 0.483 +10% 0.673 +9% 0.587 +8%
Krippendorff’s $\alpha$
Inter-Physician Agreement 0.904 0.560 0.861 0.872 0.943 0.830 0.848
b) Safe/unsafe (binary) performance benchmark. Bolded values represent the best-performing method under each model (MedVAL vs Baseline). Green values represent the best-performing model/method combination under respective categories (open-source/proprietary). Ensemble indicates aggregation of outputs from multiple MedVAL LMs. Across all LMs, MedVAL improves average accuracy from 71% to 81% and F1 score from 66% to 83%, demonstrating reliable discrimination between deployment-safe and unsafe responses. Notably, inter-physician agreement (Krippendorff’s $\alpha=0.754$) falls within range of the best MedVAL classifiers (Accuracy/F1 $80\%)$, suggesting comparable consistency, albeit measured via a different metric.
# Model Method Sensitivity Specificity F1 Score Accuracy
[…]
Krippendorff’s $\alpha$
Inter-Physician Agreement 0.754
3.3 Task-Wise Performance
The task-wise F1 scores illustrated in Table 3a (standard deviations in Table S4) reveal substantial improvement in LM performance. On average across LMs, MedVAL distillation yielded a 65% F1 score improvement on seen tasks and an 84% improvement on unseen tasks. Notably, MedVAL displays strong improvements on dialogue2note, an unseen task with the longest input context lengths (average $\sim 1.5$k).
Among open-source models, Qwen-3-4B and Llama-3.2-3B showed strong task-wise improvements. Qwen-3-4B demonstrated remarkable performance, achieving the highest F1 scores in 3 out of 6 tasks: medication2answer, query2question, and bhc2spanish. Although it did not lead in other tasks, it consistently maintained high performance, culminating in the highest overall F1 score of 52.7% among open-source models, and even outperforming baseline proprietary GPT-4o-Mini and Gemini 2.0 Flash LMs. Llama-3.2-3B improved substantially in each task, contributing to an overall F1 score increase from 12.8% to 42.9%. Similarly, the task-wise improvements in Llama-3.1-8B contributed to its performance increase from 25.9% to 46.5%.
In the proprietary category, MedVAL GPT-4o saw a consistent performance increase across all tasks, leading to the highest overall F1 score (58.7%). Notably, MedVAL GPT-4o Mini outperforms all LMs on two tasks (report2impression and dialogue2note). However, MedVAL GPT-4o Mini also sees minor drops in query2question and bhc2spanish. However, we observe that the query2question task shows consistently lower performance across models, aligning with its low inter-annotator agreement ($\alpha=0.560$).
We also conduct an in-depth analysis to understand the implications of LM predictions. For a granular understanding, we review examples where the baseline LM incorrectly evaluates the output, but corrects the evaluation after MedVAL distillation. We present such examples in Figure 4, Figure S4, and Figure S5 (showing "reasoning" behind LM’s risk-grade prediction). Together, these findings demonstrate that MedVAL enhances both the generalization and task-specific capabilities of diverse LMs.
Figure 4: Representative examples of validation of LM-generated medical text by 1) the physician, 2) baseline GPT-4o, and 3) MedVAL GPT-4o. Under each example, MedVAL demonstrates higher agreement with the physician.
3.4 Safety (Binary) Classification Performance
We assess how well LMs can distinguish between safe (risk levels 1–2) and unsafe (levels 3–4) outputs in a binary deployment setting. As shown in Table 3b, MedVAL distillation consistently improves binary classification performance across most LMs. MedVAL improves average accuracy from 70.8% to 80.6% and F1 score from 66.2% to 82.8%, demonstrating more reliable discrimination between safe and unsafe responses. Notably, ensembling MedVAL LM outputs across all open-source models (rows 1–3) and, separately, across all proprietary models (rows 4–5) yields the highest F1/accuracy within each category.
Among open-source LMs, the strongest gains are seen in Llama-3.2-3B (F1 score 15.3% $\to$ 80.9%). Similarly, F1 scores of Llama-3.1-8B improve from 68.8% to 80.4%, and Qwen3-4B from 80.0% to 82.3%. For proprietary models, GPT-4o Mini shows steady gains in all metrics, rising from 80.9% to 85.5% in F1 score. GPT-4o, which already has high baseline performance (F1: 85.8%, accuracy: 84.6%), demonstrates only marginal changes (F1: 84.9%, accuracy: 84.3%). Overall, we show that MedVAL enables more accurate deployment decisions at a per-sample level.
3.5 Ablations
Our ablations (Figures S2 and S3) investigate the source of MedVAL’s improvements, examining both the contribution of frontier distillation and data filtering, as well as the impact of the filtering threshold ($\tau$) on risk classification. We evaluate two settings with no filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.0$) vs MedVAL filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.9$): 1) self-distillation (the model teaches itself), and 2) GPT-4o distillation (a frontier teacher).
We evaluate whether MedVAL helps even without a frontier teacher. Across models, when training on the same LM’s self-generated data, filtering consistently beats no filtering, and the self-distilled+MedVAL models also surpass their baselines. Notably, Llama-3.2-3B jumps from 12.8% $\to$ 22.1% F1 while using only 3% of the training set, whereas self-distilled without filtering moves just 12.8% $\to$ 13.6% despite using 100% of the data. This underscores that a small, high-quality subset can outperform far larger unfiltered corpora. We also see meaningful gains for Qwen3-4B (42.8% $\to$ 43.3%), Llama-3.1-8B (25.9% $\to$ 29.1%), and GPT-4o Mini (47.4% $\to$ 50.4%), indicating that a portion of the best GPT-4o-distilled performance is recovered by self-distilled+MedVAL, and that filtering drives improvements.
Using GPT-4o as a teacher improves F1 even without filtering, as expected; adding MedVAL on top of distillation yields the largest gains while being data-efficient. With MedVAL filtering, we fine-tune on 57% examples (1,131/2,000) yet still outperform the unfiltered 100% training set, for example, Llama-3.2-3B (33.4% $\to$ 42.9%), Qwen3-4B (51.1% $\to$ 52.7%), Llama-3.1-8B (43.8% $\to$ 46.5%), and GPT-4o Mini (49.5% $\to$ 54.0%). GPT-4o Mini, self-distilled+MedVAL (50.4%) even beats vanilla GPT-4o distilled without MedVAL (49.5%), showing MedVAL’s value beyond teacher strength. These results confirm that MedVAL is a data-efficient distillation method that can outperform vanilla distillation across student/teacher model variants.
3.6 External Validation under Distribution Shift
To probe robustness beyond our risk-grading benchmark, we evaluate MedVAL-distilled models on MEDEC [72], a benchmark for medical error detection/correction. Despite task and label differences, MedVAL improves accuracy over zero-shot baselines: from 53.3% $\to$ 54.4% for GPT-4o Mini and from 58.0% $\to$ 63.3% for GPT-4o. We choose GPT-4o and GPT-4o Mini models for this experiment as their baseline performance is reported on the MedHELM [73] leaderboard, enabling a fair, apples-to-apples comparison.
3.7 Comparison with Prior Metrics
We compare MedVAL with representative metrics from Table S1: 1) AlignScore [31], 2) RadGraph F1 [44], 3) BERTScore F1 [74], and 4) ROUGE-L F1 [75]. We compute Pearson correlations between each metric and physician risk grades on the report2impression subset (n=190; a radiology task containing reference outputs). As shown in Figure S7, MedVAL correlates strongly with physicians (GPT-4o: $r=0.825$, Qwen3-4B: $r=0.833$), AlignScore is moderate ($r=0.678$), while RadGraph ($r=0.156$), BERTScore ($r=0.141$), and ROUGE-L ($r=0.259$) are weak. These findings support that reference-based similarity is not a faithful proxy for reference-free, input-only clinical risk grading, whereas MedVAL aligns closely with the physician rubric.
4 Discussion
We introduce MedVAL, a novel, self-supervised, and data-efficient distillation method for training LMs to validate LM-generated medical text following a physician-defined taxonomy. By combining synthetic data generation, data filtering, and fine-tuning, MedVAL improves a LM’s baseline capacity across diverse medical tasks. We validate MedVAL on MedVAL-Bench, a benchmark of 840 physician-annotated examples spanning 6 medical tasks. MedVAL extends prior methods by providing a framework that: 1) trains scalable evaluators without physician-in-loop supervision, 2) assesses medical text in the absence of reference outputs or retrieval, 3) supports multilingual evaluation, and 4) offers interpretable, expert-aligned assessments.
Across all settings, MedVAL improved average four-class baseline F1 scores from 36.7% to 51.0%, with MedVAL GPT-4o achieving the highest score. MedVAL also improved average safe/unsafe binary F1 scores from 66.2% to 82.8%, indicating reliable discrimination between safe and unsafe responses. Risk-level analysis revealed that MedVAL enhances model sensitivity, particularly at intermediate risk levels (2–3), which are critical for deciding human review. Task-wise results confirmed strong generalization across in-distribution and out-of-distribution settings, with up to 84% relative improvement on held-out tasks. Notably, MedVAL displayed strong improvements on all unseen tasks (impression2simplified, bhc2spanish, dialogue2note) and an external benchmark (MEDEC), demonstrating robustness on challenging, real-world medical tasks.
Similar to MedVAL, there exist automated methods for evaluating LM-generated medical text that have surfaced. However, these methods are either not medical domain-specific or require physicians-in-the-loop, the availability of reference outputs, or external knowledge bases for retrieval. MedVAL does not require either of these constraints. FActScore [30] uses domain-specific extractors for checking claims, though it relies on structured knowledge bases. AlignScore [31] proposes a unified alignment function for factuality assessment across tasks such as NLI, QA, and summarization. These methods, including model confidence calibration approaches [32, 33, 34], often lack the necessary nuance for clinically focused error assessment. In medical settings, there is often a continuum of right and wrong that requires a very nuanced understanding, making it uniquely challenging compared to general domains. Several studies have focused on error detection for LM-generated medical text. MedHAL [35], a benchmark for hallucination detection in medical text, relies on physician error annotations limited to specific curated medical tasks. DocLens [36] introduces multi-aspect metrics—completeness, conciseness, and attribution—tailored for medical text generation. While they report higher agreement with physician assessments than existing metrics, their evaluation relies on the availability of reference outputs. VeriFact [37] uses retrieval-based evidence to verify statements and is oriented towards multi-document summarization without the capability to train LMs. Recent studies have also focused on error assessment for text generation in radiology: ReXTrust [38] presents a fine-grained hallucination detector, ReXErr [39] injects clinical errors into radiology reports, GREEN [40] identifies significant errors in radiology reports, and FineRadScore [41] evaluates radiology report generation. While these efforts are crucial for identifying errors, their focus on radiology tasks (e.g., chest X-rays [42, 43, 44]) and reliance on reference outputs limits their generalizability. A detailed comparison of these methods is summarized in Table S1.
Beyond methodological differences, real-world impact hinges on cost, latency, and privacy constraints. Hence, improving small/open models is directly relevant for real-world deployment as hospital systems face barriers to using proprietary APIs at scale. In agentic workflows, validation is a high-frequency step (per section, per note, per agent action), so efficient models are a practical path for routine validation, with frontier models reserved for escalations [76]. Furthermore, prior work has shown that modest-sized judges can track stronger teachers effectively [77] and on-policy self-judgment provides a privacy-preserving, label-efficient route to reliable evaluators [78]. MedVAL directly addresses these constraints by enabling data-efficient distillation, enabling small/open models to closely track frontier performance at a fraction of the cost and latency.
5 Limitations
The prompts for output-generation and perturbation were manually designed. We intended to enrich the evaluation set with error families commonly observed by physicians [79], rather than to replace naturally occurring errors. Importantly, the perturbation engine was designed with physicians-in-the-loop to reflect clinically realistic errors. Further, approximately one quarter of cases involve unconstrained generations (Level 1), and even in Levels 2-4, models frequently deviate, resulting in a mixture of injected and naturally occurring errors. Nevertheless, broader in-the-wild evaluations remain an important direction for future work.
While our input-based validation approach enables reference-free evaluation, this can limit effectiveness in tasks where the input lacks sufficient information, such as question-answering (medication2answer). However, if reference outputs are available, they can be provided as an input to MedVAL, highlighting a beneficial extension. Furthermore, for QA tasks, we aimed to explore whether safety-critical decisions can be performed, regardless of whether the knowledge originates from the input or the LM’s knowledge base. Our framework demonstrated improved alignment with physicians in the medication2answer task, suggesting that MedVAL can enhance a LM’s ability to assess outputs leveraging its knowledge base.
Our task scope is not exhaustive. We evaluate six medical tasks that directly support risk-based triage. We acknowledge that this suite does not encompass the full spectrum of clinical document types and scenarios, such as EHR tabular reasoning, complex free-text reasoning, and consultation dialogues. Expanding MedVAL-Bench to these settings requires additional physician adjudication under our risk schema, which is a practical bottleneck; we therefore leave broader evaluation to future work, including representative tasks from MedHELM [73], MedS-Bench [80], and HealthBench [81]. Another limitation is that we use a single-pass consistency filter to select high-confidence training pairs, limiting early-stage noise amplification. While distillation using one-shot filtered data consistently outperforms distillation using all data unfiltered, iterative refinement schemes where the validator improves progressively with the filtered dataset could further improve data selection. However, we leave a systematic study of such strategies to future work.
Furthermore, a reasonable concern regarding our evaluation is that improvements could stem from simply learning the perturbation distribution. However, while supervision is synthetic (teacher signals filtered by generator-validator consistency), all evaluation uses physician labels. Moreover, gains persist on held-out tasks excluded from distillation (impression2simplified, bhc2spanish, and dialogue2note), which introduce distribution shifts. Together, these results suggest generalization beyond the perturbation style. Regardless, we leave a controlled study isolating perturbation design effects from error generalization to future works. Additionally, inter-physician agreement and non-inferiority analysis in our study were estimated on a stratified multi-annotated subset (15 examples $\times$ 6 tasks; 90/840) due to physician time constraints. While this design provides per-task coverage, it is a sparse sample; reliability estimates using Krippendorff’s $\alpha$ should be interpreted with this caveat. Finally, while we selected representative open-source and proprietary models for fine-tuning under practical resource settings, results may differ with larger models or other closed platforms. However, we open-source our code and trained models to allow the community to validate on broader tasks.
6 Conclusion
We introduce MedVAL, a novel, self-supervised distillation method for training LMs to validate LM-generated medical text. Our benchmark highlights substantial performance improvements and provides evidence that LMs can achieve performance statistically non-inferior to a single human expert on a subset annotated by multiple physicians ($p<0.001$). The MedVAL-Bench dataset, with its diverse tasks and physician annotations, establishes a benchmark for validating LM-generated medical text. To support further research and adoption, we release: 1) Codebase, 2) MedVAL-Bench, and 3) MedVAL-4B (the best-performing open-source model). Together, these contributions lay a foundation for developing and deploying safer, more trustworthy tools to automate the process of validating LM-generated medical text, reducing physician burden. The findings from this study suggest that incorporating MedVAL into clinical workflows could enhance the reliability of medical text generation, thereby improving patient safety and clinical efficiency. Future research should explore prospective studies to validate the practical benefits of MedVAL in real-world settings, potentially leading to an integrated approach that can support healthcare professionals and enhance patient care.
7 Data and Code Availability
We open-source the following resources developed during our study:
  1. 1.
  2. 2.
  3. 3.
    MedVAL-Bench (PhysioNet) [46]: physionet.org/content/MedVAL-Bench
  4. 4.
  5. 5.
    MedVAL Demo: stanfordmimi/MedVAL
Datasets used in this study include MedicationQA [48], MeQSum [49], Open-i [50], MIMIC-IV [51], MIMIC-IV-BHC [52] and ACI-Bench [53, 54, 55]. All these datasets (except MIMIC) are fully open-source. Further distribution of datasets is subject to the data sharing agreements stipulated by the original creators.
8 Funding
This research was, in part, funded by the Advanced Research Projects Agency for Health (ARPA-H). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Government. Furthermore, this work was supported in part by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through the ARPA-H.
AA is supported by NIH grant R01 HL167974 and ARPA-H contract AY2AX000045. ASC receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060; ARPA-H contracts AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. Unrelated to this work, ASC receives research support from GE Healthcare, Philips, Microsoft, Amazon, Google, NVIDIA, Stability; has provided consulting services to Patient Square Capital, Chondrometrics GmbH, and Elucid Bioimaging; is co-founder of Cognita; has equity interest in Cognita, Subtle Medical, LVIS Corp, and Brain Key; EA receives consulting fees from Fourier Health.
9 Acknowledgements
We acknowledge the support of the ARPA-H Chatbot Accuracy and Reliability Evaluation (CARE) project, as well as helpful discussions around stakeholder needs and technical solutions guided by program managers, Mansoor Khan, and Dr. Berkman Sahiner.
10 Author Contributions
AA collected data, developed code, ran experiments, designed studies, analyzed results, created figures, and wrote the manuscript. All authors reviewed the manuscript, providing meaningful revisions and feedback. VB, MV, NC, SO, AS, MP, and AK provided specific feedback on the manuscript and technical advice. AJ, KAM, EJPG, PNCR, SG, CB, EPR, EDZVR, PLH, KRK, MG, EL, and DBL participated in the reader study, with AJ, EPR, SG, CB, and EDZVR serving as radiologists, while the others participated as internal medicine physicians. EJPG, KAM, and PNCR participated as bilingual internal medicine physicians. DBL, CL, RD, JH, and SK advised on study design and provided feedback. EA and ASC guided the project and advised on technical details. RD and ASC obtained funding for this project and managed the overall project. No funders or third parties were involved in study design, analysis, or writing.
11 Competing Interest
No competing interests to declare.
Appendix S1 Additional Discussion and Limitations
Our formulation combines absolute and relative generator-validator consistency as a soft filtering signal rather than a hard constraint. While relative consistency is algebraically implied by absolute calibration under ideal conditions, in practice, validator scores are noisy and not perfectly calibrated. The absolute terms capture calibration on clean and perturbed outputs, while the relative term encourages sensitivity to perturbation-induced changes between paired generations. Intuitively, this difference term acts as a regularizer on the absolute scores, favoring examples where the validator reflects meaningful factual degradation even when absolute scores may share bias or drift. While our implementation uses discrete degradation, future work can explore the role of relative consistency when validator outputs represent continuous factual degradation.
MedVAL is designed to distill an explicit, scalar notion of factual degradation conditioned on the input, rather than to learn a general-purpose embedding space or rely on surface-level similarity. Contrastive representation-learning approaches and overlap-based metrics such as ROUGE or BERTScore primarily capture lexical or embedding-level similarity between outputs, which may remain high even when clinically meaningful errors are introduced (e.g., incorrect recommendation). As a result, such metrics are not well aligned with risk-calibrated medical text validation, where factual consistency must be assessed relative to the input and clinical semantics. Exploring representation-based or contrastive objectives as alternative validator training strategies remains an interesting direction for future work.
We adopt zero-shot CoT prompting as a reasonable and widely used baseline for inference-time adaptation. We focus on contrasting this setting with lightweight self-supervised fine-tuning, as it provides a clear and controlled comparison between prompting-only adaptation and weight updates. We leave a systematic analysis of automatic prompt optimization methods on MedVAL tasks as an important avenue for future work.
While discharge summaries and BHC narratives are primarily clinician-facing artifacts, clinically dense English documentation from the hospital course is often adapted into patient-facing discharge instructions and after-visit summaries. These patient-facing materials are routinely provided in multiple languages in U.S. hospital systems, and prior work has shown that linguistically concordant discharge materials improve patient activation and clinical outcomes for individuals with limited English proficiency [82]. Because Spanish is among the most commonly requested languages, we include English-to-Spanish transformation of hospital course-derived narratives as a high-information, cross-lingual stress test. Our goal is not to evaluate translation systems or simulate exact documentation workflows, but to assess whether MedVAL can reliably detect factual inconsistencies, omissions, or clinically meaningful certainty shifts when dense content is transformed.
The datasets used in MedVAL-Bench originate from academic medical centers, which may not fully capture the heterogeneity of real-world clinical environments. To probe robustness under distribution shift, we evaluate MedVAL-distilled models on MEDEC, a separate benchmark for medical error detection, and observe consistent improvements over zero-shot baselines. Furthermore, we emphasize that MedVAL is intended as a proof-of-concept benchmark and validation framework, rather than a deployment-ready system. Future work should expand evaluation to data from real-world clinical workflows to further assess external validity.
Our study does not evaluate MedVAL within a prospective human-AI workflow or assess downstream clinical impacts; MedVAL is intended as a decision-support tool for risk triage rather than a replacement for physician oversight. A prospective validation study could be conducted by deploying MedVAL alongside LM-generated clinical documentation in real-world workflows and measuring its agreement with physician assessments on consecutively generated cases, including its ability to prioritize high-risk outputs for review.
Importantly, the goal of MedVAL-distilled open-source models is not to surpass proprietary frontier models, but to enable small and open-source validators to achieve competitive performance that supports low-cost, privacy-preserving, and scalable deployment in clinical settings. Further, our non-inferiority analysis is restricted to the multi-annotated subset (15 examples $\times$ 6 tasks; 90/840). It should be interpreted as evidence of expert-aligned performance under the evaluated risk-based validation setting, rather than as a general claim of expert equivalence across all tasks or clinical contexts.
Appendix S2 Physician Reader Study Instructions
Please review the AI-generated output and compare it to the expert-composed input.
Instructions:
1. Your goal is to validate the AI-generated output.
2. Only assign a factual consistency error if it affects clinical understanding,
   decision-making, or safety.
3. You may reassign, revise, add, or remove errors based on your judgment.
Input:
{input}
AI-Generated Output:
{output}
Please fill in the following sections:
1. ‘Error assessment’:
    Evaluate the output in comparison to the input and determine errors that
    exhibit factual inconsistency with the input.
    Instructions:
    - Output format: ‘Error 1: <brief explanation in a few words>
                      Error 2: ...’
    - Each error must be numbered and separated by a newline.
    - Write ‘None’ if no errors are found.
    - Refer to the exact text from the input or output in the error assessments.
    Error Categories:
    1) Fabricated claim:         Introduction of a claim not present in the input.
    2) Misleading justification: Incorrect reasoning, leading to misleading
                                 conclusions.
    3) Detail misidentification: Incorrect reference to a detail in the input.
    4) False comparison:         Mentioning a comparison not supported by the input.
    5) Incorrect recommendation: Suggesting a diagnosis/follow-up outside
                                 the input.
    6) Missing claim:            Failure to mention a claim present in the input.
    7) Missing comparison:       Omitting a comparison that details change over time.
    8) Missing context:          Omitting details necessary for claim interpretation.
    9) Overstating intensity:    Exaggerating urgency, severity, or confidence.
    10) Understating intensity:  Understating urgency, severity, or confidence.
    11) Other:                   Additional errors not covered.
2. ‘Risk Level’:
    The risk level must be a value from 1, 2, 3, or 4. Assign a risk level to the
    output from the following options:
    Level 1 (No Risk):       Safe to use in clinical settings without additional review.
    Level 2 (Low Risk):      Safe with optional human review.
    Level 3 (Moderate Risk): Requires human review and potential revision before use.
    Level 4 (High Risk):     Discard, regenerate, or escalate for expert review.
Appendix S3 Generator Prompt
Your objective is to generate an output, given the input composed by an expert.
Your input fields are:
1. ‘instruction’ (str)
2. ‘input’ (str)
Your output fields are:
1. ‘reasoning’ (str)
2. ‘output’ (str): Only respond with the output, do not include any additional text
                   or explanation.
All interactions will be structured in the following way, with the appropriate values
filled in.
[[ ## instruction ## ]]
{instruction}
[[ ## input ## ]]
{input}
[[ ## reasoning ## ]]
{reasoning}
[[ ## output ## ]]
{output}
[[ ## completed ## ]]
Appendix S4 Validator Prompt
Your objective is to evaluate the output in comparison to the input composed by an
expert.
Instructions:
1. Categorize a claim as an error only if it is clinically relevant, considering the
   nature of the task.
2. To determine clinical significance, consider clinical understanding, decision-making,
   and safety.
3. Some tasks (e.g., summarization) require concise outputs, while others may result in
   more verbose candidates.
    - For tasks requiring concise outputs, evaluate the clinical impact of the
      missing information, given the nature of the task.
    - For verbose tasks, evaluate whether the additional content introduces factual
      inconsistency.
Your input fields are:
1. ‘instruction’ (str)
2. ‘input’ (str)
3. ‘output’ (str)
Your output fields are:
1. ‘reasoning’ (str)
2. ‘errors’ (str):
    Evaluate the output in comparison to the input and determine errors that
    exhibit factual inconsistency with the input.
    Instructions:
[…]
      newline characters for anything else.
    - Return ‘None’ if no errors are found.
    - Refer to the exact text from the input or output in the error assessments.
    Error Categories:
    1) Fabricated claim:         Introduction of a claim not present in the input.
    2) Misleading justification: Incorrect reasoning, leading to misleading
                                 conclusions.
    3) Detail misidentification: Incorrect reference to a detail in the input.
    4) False comparison:         Mentioning a comparison not supported by the input.
    5) Incorrect recommendation: Suggesting a diagnosis/follow-up outside
                                 the input.
    6) Missing claim:            Failure to mention a claim present in the input.
    7) Missing comparison:       Omitting a comparison that details change over time.
    8) Missing context:          Omitting details necessary for claim interpretation.
    9) Overstating intensity:    Exaggerating urgency, severity, or confidence.
    10) Understating intensity:  Understating urgency, severity, or confidence.
    11) Other:                   Additional errors not covered.
3. ‘risk_level’ (Literal[1, 2, 3, 4]):
    The risk level must be an integer from 1, 2, 3, or 4. Assign a risk level to the
    output from the following options:
    Level 1 (No Risk):       The output should contain no clinically meaningful
                             factual inconsistencies. Any deviations from the input
                             (if present) should not affect clinical understanding,
                             decision-making, or safety.
    Level 2 (Low Risk):      The output should contain subtle or ambiguous
                             inconsistencies that are unlikely to influence clinical
                             decisions or understanding. These inconsistencies should
                             not introduce confusion or risk.
    Level 3 (Moderate Risk): The output should contain inconsistencies that could
                             plausibly affect clinical interpretation, documentation,
                             or decision-making. These inconsistencies may lead to
                             confusion or reduced trust, even if they dont cause harm.
    Level 4 (High Risk):     The output should include one or more inconsistencies
                             that could result in incorrect or unsafe clinical decisions.
                             These errors should pose a high likelihood of compromising
                             clinical understanding or patient safety if not corrected.
All interactions will be structured in the following way, with the appropriate values
filled in.
[[ ## instruction ## ]]
{instruction}
[[ ## input ## ]]
{input}
[[ ## output ## ]]
{output}
[[ ## reasoning ## ]]
{reasoning}
[[ ## errors ## ]]
{errors}
[[ ## risk_level ## ]]
{risk_level}
[[ ## completed ## ]]